OpenStudy (anonymous):

Now it is your turn to plan the trajectories required to launch a spacecraft through a specific route in space. The launch area is identified on the map below. A coordinate plane is shown with a point at 1, 2 labeled, Launch Area. Launch Area:___(1, 2)___ Point A:___(0, 3)___ Point B:___(-3, 0)___ Point C:___(-1, -4)___ Graph the coordinates of the specifics points in space your spacecraft will travel to. You must show your work on each question below. Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. Determine the equation

OpenStudy (anonymous):

i just need Number 6!!!! my answer for number 2 is x+y=3 3- y-3=1(x-0) 4- y=-2x-6 5- y=-1x+3 6- ????????